# MGDA Converges under Generalized Smoothness, Provably

### v4

@@ A.3 Implementation details
Generalized smoothness illustration. To illustrate the relation between gradient norms and local smoothness, we run SGD on each task separately without the warm start process. Since there is no weight update process, we only need to choose $\alpha=0.0005$ for both tasks.
Appendix B Algorithm
@@
Appendix C Detailed Proofs for Average CA Distance
C.1 Formal version and proof of Theorem 1
Let $c_{1}>0,c_{2}>0$, $c_{3}\geq 0$ and $F>0$ be some constants such that
$\Delta+c_{1}+{c_{2}}+c_{3}\leq F.$ (8)
@@ C.1 Formal version and proof of Theorem 1
Suppose Assumptions 1 and 2 are satisfied, and we choose constant step sizes that $\beta\leq\frac{1}{4KM^{2}},\alpha\leq\min\left(c_{1}\beta,\frac{1}{2\ell(M+1)},\frac{1}{M\ell(M+1)}\right),T\geq\max\left(\frac{10\Delta}{\alpha\epsilon^{2}},\frac{10}{\epsilon^{2}\beta}\right)\sim\Theta(\epsilon^{-2})$, and $\rho\leq\min\left(\frac{\epsilon^{2}}{20},\sqrt{\frac{\epsilon^{2}}{10\beta}},\frac{c_{2}}{2T\alpha},\sqrt{\frac{c_{3}}{T\alpha\beta}}\right)\sim\mathcal{O}(\epsilon^{2})$. We have that
@@
Base Case: since all $c_{1},c_{2},c_{3},M$ are non-negative, according to (8) we have that $f_{i}(x_{0})-f_{i}^{*}\leq\Delta\leq F$ holds for any $i\in[K]$.
@@
where the second inequality is due to $\alpha\leq\frac{1}{M\ell(M+1)}$ and the last inequality is due to $\|\nabla f_{i}(x_{t})\|\leq M$. Based on Assumption 2, Definition 2 and Lemma 3.3 in [21], we have the following descent lemma:
@@
Plug (C.1) into (10), and we can show that
@@
Taking sums of (C.1) from $t=0$ to $k$, for any $w\in\mathcal{W}$ we have that
@@
since we have that $\frac{\alpha}{\beta}\leq c_{1},T\alpha\beta\rho^{2}\leq c_{3}$ and $2T\alpha\rho\leq c_{2}$. Now we finish the induction step and can show that $f_{i}(x_{k+1})-f_{i}^{*}\leq F$ and (C.1) hold for all $k<T$ and $i\in[K]$.
Specifically, for $\alpha<\frac{1}{2\ell(M+1)},\beta\leq\frac{1}{4KM^{2}}$, according to (C.1), for $k=T-1$ we have that
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}\leq\frac{2(F(x_{0})w-F^{*}w)}{\alpha T}+\frac{2}{\beta T}+2\beta\rho^{2}+4\rho\leq\epsilon^{2},$
@@
(Lemma 3.5 in [21]) If a function $f$ is $\ell$-smooth, we have that
@@ C.2 Proof of Corollary 1
$\|\nabla F(x_{t})w_{t}-\nabla F(x_{t})w_{t}^{*}\|^{2}\leq\|\nabla F(x_{t})w_{t}\|^{2}-\|\nabla F(x_{t})w_{t}^{*}\|^{2}\leq\|\nabla F(x_{t})w_{t}\|^{2},$
where the first inequality follows from the optimal condition. Then we have
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}-\nabla F(x_{t})w_{t}\|^{2}\leq\frac{1}{T}\sum_{t=0}^{T-1}|\|\nabla F(x_{t})w_{t}\|^{2}=\mathcal{O}(\epsilon^{2}),$
@@
C.3 Formal Version and Its Proof of Theorem 2
@@
where $\Delta=\max_{i\in[d]}\{f_{i}(x_{0})-f^{*}\}$. Let $L_{0}>0,L_{1}>0,b_{1}>0,b_{2}>0,b_{3}>0$ be some constants.
For $\alpha\leq\min\{\frac{\ell(M+1)}{2\max(1,M)},c_{1}\beta,\rho\min\left\{\frac{1}{4L_{0}^{2}\ell(M+1)^{2}},\frac{b_{1}^{2}}{(3M\sqrt{K}L_{0}+M\sqrt{K}L_{1})^{2}},\frac{b_{2}}{\ell(M+1)L_{0}^{2}}\right\},\frac{c_{2}}{\sqrt{2T}M(3\sigma+\sigma^{2}),},\\ \frac{\sqrt{c_{3}}}{\sqrt{\ell(M+1)\sigma^{2}T}},\frac{\delta\epsilon^{2}}{48\sigma^{2}\ell(M+1)}\}$,
$\beta\leq\min\{\frac{c_{4}}{4K\alpha T(M^{2}+\sigma^{2})^{2}},\frac{\delta\epsilon^{2}}{192K(M^{2}+\sigma^{2})^{2}},\frac{b_{3}\rho}{8KM^{2}L_{0}^{2}+4KL_{1}^{2}}\},\\ \rho\leq\min\{\frac{c_{5}}{\alpha T},{\frac{\sqrt{c_{6}}}{\sqrt{\alpha\beta T}}},\frac{\delta\epsilon^{2}}{48},\sqrt{\frac{\delta\epsilon^{2}}{48\beta}},\frac{\delta L_{0}^{2}}{24K\sigma^{2}\alpha T},\frac{\delta L_{1}^{2}}{8K^{2}\sigma^{4}\alpha T}\}$,
and $T\geq\max\{\frac{48\Delta+48c_{1}}{\delta\alpha\epsilon^{2}},\frac{4608M^{2}(3\sigma+\sigma^{2})^{2}}{\delta^{2}\epsilon^{4}}\}$ such that
$b_{1}+b_{2}+c_{1}+\alpha\rho(1+\beta\rho)+4\alpha\beta KM^{4}\leq\frac{F}{2}.$ (15)
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}\leq\epsilon^{2},$
@@
($\mathbb{P}(\{\tau_{3}<T\})\leq\sum_{t=0}^{T-1}\sum_{i=1}^{K}\sum_{j=1}^{K}\mathbb{P}(\|\varepsilon_{t,2,i}\|\|\varepsilon_{t,3,i}\|>\frac{L_{1}}{\sqrt{\alpha\rho}})\leq\frac{K^{2}\sigma^{4}{\alpha\rho}T}{L_{1}^{2}}\leq\frac{\delta}{8}.$ (17)
@@
According to (C.4) shown in Lemma 1, for any $i\in[K]$ and $t=\tau$ we have that
@@
$\mathbb{E}[f_{i}(x_{\tau})-f_{i}^{*}]\leq\frac{\delta}{8}.$
@@
Convergence of $\frac{1}{T}\mathbb{E}\left[\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}\Big{|}\tau=T\right]$. Based on (C.4) in Lemma 1, we have that
@@
where the last inequality is due to (16), (17), (18), and (19). This completes the proof. ∎
C.4 Proof of Lemma 1
@@
Combine (C.4) and (C.4), and we can get that
@@
Taking expectation and sum up (C.4) from $t=0$ to $\tau-1$, we have that
@@
Based on (C.4) and (C.4), we have that
@@
Appendix D Detailed Proofs for Iteration-wise CA Distance
@@
Since we set $w_{0,\rho}(x_{t})=w_{0,\rho}(x_{t+1})$ and $\|w_{Q,\rho}(x_{t+1})\|\leq 1$, telescoping the above inequality from $t=0,1,...,T-1$ gives,
@@
where $(i)$ follows from eq. 26 and $(ii)$ follows from the convergence of PGD (Theorem 1.1, [4]) on $\rho$-strongly convex objectives that
@@
$-2\langle w_{t,\rho}^{*},\nabla F(x_{t})^{\top}\nabla F(x_{t})w_{t}^{*}\rangle\leq-2\|\nabla Fw_{t}^{*}\|^{2}.$
@@
$\|R(x_{t})\|\leq\alpha^{2}\frac{\ell(M+1)}{2}\|\nabla F(x_{t})w_{t}\|^{2}.$
@@ D.1 Proof of Theorem 3
Suppose Assumptions 1 and 2 are satisfied. We choose $\beta^{\prime}\leq\frac{1}{M^{2}},N\sim\mathcal{O}(\epsilon^{-2}),\beta\leq\min\left(\frac{\epsilon^{2}\rho}{C_{1}^{2}},\epsilon^{2}\right),\alpha\leq\min\left(c_{1}\beta,\frac{1}{\ell(M+1)},\frac{\beta\rho\epsilon}{2L_{w}\sqrt{M}},\frac{\rho\epsilon^{2}}{2L_{w}MC_{1}}\right),T\geq\max\left(\frac{10\Delta}{\alpha\epsilon^{2}},\frac{10}{\epsilon^{2}\beta}\right)\sim\Theta(\epsilon^{-11})$, and $\rho\leq\min\left(\frac{\epsilon^{2}}{20},\frac{c_{2}}{2T\alpha}\right)\sim\mathcal{O}(\epsilon^{2})$. The CA distance in every iteration takes the order of $\mathcal{O}(\epsilon)$.
@@
where $(i)$ follows from the update rule in Algorithm 1, Lemma 3, and Young’s inequality. Then substituting section D.1, eq. 31 and section D.1 into eq. 29, we have
@@
where the last inequality follows from Lemma 5 and $C_{1}=\sqrt{K}M^{2}+\rho$. Then we do telescoping over $t=0,1,...,T-1$
@@
Then recalling that $L_{w}=\mathcal{O}(\frac{1}{\rho})$ and substituting the above inequality into section D.1, we have
@@
Since we run projected gradient descent for the strongly convex function $J(w_{n})=\frac{1}{2}\|\nabla F(x_{0})w_{n}\|^{2}+\frac{\rho}{2}\|w_{n}\|^{2}$ in the N-loop in Algorithm 1, according to Theorem 10.5 [14], we have by choosing $\beta^{\prime}\in(0,\frac{1}{M^{2}}]$
@@
Thus, $\|w_{0}-w_{0,\rho}^{*}\|=\mathcal{O}(\epsilon)$ as $N\sim\mathcal{O}(\rho^{-1})$. CA distance takes the order of $\epsilon$ in every iteration by choosing $\rho\sim\mathcal{O}(\epsilon^{2}),\beta\sim\mathcal{O}(\epsilon^{4})$, $\alpha\sim\mathcal{O}(\epsilon^{9})$, and $N\sim\mathcal{O}(\epsilon^{-2})$. The proof is complete. ∎
D.2 Formal Version and Its Proof of Theorem 4

### v5

@@ B.4 Implementation details
Generalized smoothness illustration. To illustrate the relation between gradient norms and local smoothness, we follow the method in Section H.3 in Zhang et al. (2019)and run SGD on each task separately without the warm start process. The code is available in https://github.com/JingzhaoZhang/why-clipping-accelerates. Since there is no weight update process, we only need to choose $\alpha=0.0005$ for both tasks.
Appendix C Algorithm
@@
Appendix D Detailed Proofs for Average CA Distance
D.1 Formal version and proof of Theorem 1
Let $F>0$ be some constants such that
$\Delta+{3}\leq F.$ (7)
@@ D.1 Formal version and proof of Theorem 1
Suppose Assumptions 1 and 2 are satisfied, and we choose constant step sizes that $\beta\leq\frac{1}{4KM^{2}},\alpha\leq\min\Big{\{}{\beta},\frac{1}{2\ell(M+1)},\frac{1}{M\ell(M+1)}\Big{\}},T\geq\max\Big{\{}(\frac{10\Delta}{\alpha\epsilon^{2}},\frac{10}{\epsilon^{2}\beta}\Big{\}})=\Theta(\epsilon^{-2})$, and $\rho\leq\min\left(\frac{\epsilon^{2}}{20},\sqrt{\frac{\epsilon^{2}}{10\beta}},{\frac{1}{2T\alpha},\sqrt{\frac{1}{T\alpha\beta}}}\right)=\mathcal{O}(\epsilon^{2})$. We have that
@@
Base Case: since $M$ is non-negative, according to equation 7 we have that $f_{i}(x_{0})-f_{i}^{*}\leq\Delta\leq F$ holds for any $i\in[K]$.
@@
where the second inequality is due to $\alpha\leq\frac{1}{M\ell(M+1)}$ and the last inequality is due to $\|\nabla f_{i}(x_{t})\|\leq M$. Based on Assumption 2, Definition 2 and Lemma 3.3 in (Li et al., 2024a), we have the following descent lemma:
@@
Plug equation D.1 into equation 9, and we can show that
@@
Taking sums of equation D.1 from $t=0$ to $k$, for any $w\in\mathcal{W}$ we have that
@@
since we have that $\frac{\alpha}{\beta}\leq 1,T\alpha\beta\rho^{2}\leq 1$ and $2T\alpha\rho\leq 1$. Now we finish the induction step and can show that $f_{i}(x_{k+1})-f_{i}^{*}\leq F$ and equation D.1 hold for all $k<T$ and $i\in[K]$.
Specifically, for $\alpha<\frac{1}{2\ell(M+1)},\beta\leq\frac{1}{4KM^{2}}$, according to equation D.1, for $k=T-1$ we have that
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}\leq\frac{2(F(x_{0})w-F^{*}w)}{\alpha T}+\frac{2}{\beta T}+2\beta\rho^{2}+4\rho=\mathcal{O}(\epsilon^{2}),$
@@
(Lemma 3.5 in Li et al. (2024a)) If a function $f$ is $\ell$-smooth, we have that
@@ D.2 Proof of Corollary 1
$\|\nabla F(x_{t})w_{t}-\nabla F(x_{t})w_{t}^{*}\|^{2}$
$=$ $\|\nabla F(x_{t})w_{t}\|^{2}+\|\nabla F(x_{t})w_{t}^{*}\|^{2}-2\langle\nabla F(x_{t})w_{t},\nabla F(x_{t})w_{t}^{*}\rangle$
$\leq$ $\|\nabla F(x_{t})w_{t}\|^{2}+\|\nabla F(x_{t})w_{t}^{*}\|^{2}-2\|\nabla F(x_{t})w_{t}^{*}\|^{2}$
$=$ $\|\nabla F(x_{t})w_{t}\|^{2}-\|\nabla F(x_{t})w_{t}^{*}\|^{2}$
$\leq$ $\|\nabla F(x_{t})w_{t}\|^{2},$
where the first inequality holds because of the optimality condition of $w_{t}^{*}=\arg\min_{w\in\mathcal{W}}\frac{1}{2}\|\nabla F(x_{t})w\|^{2}$ that for any $w\in\mathcal{W}$,
$\langle w,\nabla F(x_{t})^{\top}\nabla F(x_{t})w_{t}^{*}\rangle\geq\langle w_{t}^{*},\nabla F(x_{t})^{\top}\nabla F(x_{t})w_{t}^{*}\rangle=\|\nabla F(x_{t})w_{t}^{*}\|^{2}.$
Then we have
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}-\nabla F(x_{t})w_{t}^{*}\|^{2}\leq\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}=\mathcal{O}(\epsilon^{2}),$
@@
D.3 Formal Version and Its Proof of Theorem 2
@@
where $\Delta=\max_{i\in[d]}\{f_{i}(x_{0})-f^{*}\}$. Let $L_{0}>0,L_{1}>0,b_{1}>0,b_{2}>0,b_{3}>0$ be some constants, for $B_{1}=\min\left\{\frac{1}{4L_{0}^{2}\ell(M+1)^{2}},\frac{b_{1}^{2}}{(3M\sqrt{K}L_{0}+M\sqrt{K}L_{1})^{2}},\frac{b_{2}}{\ell(M+1)L_{0}^{2}}\right\}$.
$\rho\leq\min\{\frac{\delta\epsilon^{2}}{48},\sqrt{\frac{\delta\epsilon^{2}}{48}},\frac{L_{0}^{2}\delta^{2}\epsilon^{2}}{1152K\sigma^{2}(\Delta+c_{1})},\frac{L_{1}^{2}\delta^{2}\epsilon^{2}}{384K\sigma^{2}(\Delta+c_{1})}\}=\mathcal{O}(\delta^{2}\epsilon^{2})$,
$\beta\leq\min\{\frac{\delta\epsilon^{2}}{192K(M^{2}+\sigma^{2})^{2}},\frac{b_{3}\rho}{8KM^{2}L_{0}^{2}+4KL_{1}^{2}},1\}=\mathcal{O}(\min\{\frac{\delta\epsilon^{2}}{M^{4}},\frac{\rho}{M^{2}}\})$,
$\alpha\leq\min\{\frac{\ell(M+1)}{2\max(1,M)},c_{1}\beta,\rho B_{1},\frac{c_{2}}{\sqrt{2T}M(3\sigma+\sigma^{2}),},\frac{\sqrt{c_{3}}}{\sqrt{\ell(M+1)\sigma^{2}T}},\frac{\delta\epsilon^{2}}{48\sigma^{2}\ell(M+1)},\\ \frac{1}{\rho T}\min\{c_{5},\frac{\delta L_{0}^{2}}{24K\sigma^{2}},\frac{\delta L_{1}^{2}}{8K\sigma^{4}}\},\frac{c_{4}}{4K\beta T(M^{2}+\sigma^{2})^{2}},\frac{c_{6}}{\rho\beta T}\}=\mathcal{O}(\min\{\beta,\frac{\rho}{\ell(M+1)^{2}},\frac{\rho}{M^{2}},\frac{\delta}{\rho T},\frac{1}{\beta TM^{2}}\})$
and $T\geq\max\{\frac{48\Delta+48c_{1}}{\delta\alpha\epsilon^{2}},\frac{4608M^{2}(3\sigma+\sigma^{2})^{2}}{\delta^{2}\epsilon^{4}}\}=\mathcal{O}(\max\{\frac{1}{\delta\alpha\epsilon^{2}},\frac{M^{2}}{\delta^{2}\epsilon^{4}}\})$ such that
$b_{1}+b_{2}+b_{3}+c_{1}+\alpha\rho(1+\beta\rho)+4\alpha\beta KM^{4}\leq\frac{F}{2}.$ (14)
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}=\mathcal{O}(\epsilon^{2}),$
@@
($\mathbb{P}(\{\tau_{3}<T\})\leq\sum_{t=0}^{T-1}\sum_{i=1}^{K}\sum_{j=1}^{K}\mathbb{P}(\|\varepsilon_{t,2,i}\|\|\varepsilon_{t,3,j}\|>\frac{L_{1}}{\sqrt{\alpha\rho}})\leq\frac{K^{2}\sigma^{4}{\alpha\rho}T}{L_{1}^{2}}\leq\frac{\delta}{8}.$ (16)
@@
According to equation D.4 shown in Lemma 1, for any $i\in[K]$ and $t=\tau$ we have that
@@
$\mathbb{E}[f_{i}(x_{\tau})-f_{i}^{*}]\leq\frac{\delta}{8}F.$
@@
Convergence of $\frac{1}{T}\mathbb{E}\left[\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}\Big{|}\tau=T\right]$. Based on equation D.4 in Lemma 1, we have that
@@
where the last inequality is due to equation 15, equation 16, equation 17, and equation 18. This completes the proof. ∎
D.4 Proof of Lemma 1
@@
Combine equation D.4 and equation D.4, and we can get that
@@
Taking expectation and sum up equation D.4 from $t=0$ to $\tau-1$, we have that
@@
Based on equation D.4 and equation D.4, we have that
@@
Appendix E Detailed Proofs for Iteration-wise CA Distance
@@
Since we set $w_{0,\rho}(x_{t})=w_{0,\rho}(x_{t+1})$ and $\|w_{Q,\rho}(x_{t+1})\|\leq 1$, telescoping the above inequality over $Q$ gives,
@@
where $(i)$ follows from eq. 25 and $(ii)$ follows from the convergence of PGD (Theorem 1.1, (Beck & Teboulle, 2009)) on $\rho$-strongly convex objectives that
@@
$-2\langle w_{t,\rho}^{*},\nabla F(x_{t})^{\top}\nabla F(x_{t})w_{t}^{*}\rangle\leq-2\|\nabla F(x_{t})w_{t}^{*}\|^{2}.$
@@
$\|R(x_{t})\|\leq\alpha^{2}\frac{\ell(M+1)}{2}\|\nabla F(x_{t})w_{t}\|^{2}\leq\frac{\alpha^{2}K\ell(M+1)M^{2}}{2}.$
@@ E.1 Proof of Theorem 3
Suppose Assumptions 1 and 2 are satisfied. We choose $\beta^{\prime}\leq\frac{1}{M^{2}},N=\mathcal{O}(\epsilon^{-2}),{C_{1}\geq\sqrt{K}M^{2}+\rho},\beta\leq\min\left(\frac{\epsilon^{2}\rho}{C_{1}^{2}},\epsilon^{2},\frac{1}{4KM^{2}}\right),\alpha\leq\min\left(\beta,\frac{1}{\ell(M+1)},\frac{\beta\rho\epsilon}{2L_{w}\sqrt{M}},\frac{\rho\epsilon^{2}}{2L_{w}MC_{1}}\right),T\geq\max\left(\frac{10\Delta}{\alpha\epsilon^{2}},\frac{10}{\epsilon^{2}\beta}\right)=\Theta(\epsilon^{-11})$, and $\rho\leq\min\left(\frac{\epsilon^{2}}{20},\frac{1}{2T\alpha},\sqrt{\frac{\epsilon^{2}}{10\beta}},\sqrt{\frac{1}{T\alpha\beta}}\right)=\mathcal{O}(\epsilon^{2})$. The CA distance in every iteration takes the order of $\mathcal{O}(\epsilon)$.
@@
where $(i)$ follows from the update rule in Algorithm 1, Lemma 3, and Young’s inequality. Then substituting section E.1, eq. 30 and section E.1 into eq. 28, we have
@@
where the last inequality follows from Lemma 5 and $C_{1}\geq\sqrt{K}M^{2}+\rho$. Then we do telescoping over $t=0,1,...,T-1$
@@
Then recalling that $L_{w}=\mathcal{O}(\frac{1}{\rho})$ and substituting the above inequality into section E.1, we have
@@
Since we run projected gradient descent for the strongly convex function $J(w_{n})=\frac{1}{2}\|\nabla F(x_{0})w_{n}\|^{2}+\frac{\rho}{2}\|w_{n}\|^{2}$ in the N-loop in Algorithm 1, according to Theorem 10.5 (Garrigos & Gower, 2023), we have by choosing $\beta^{\prime}\in(0,\frac{1}{M^{2}}]$
@@
Thus, $\|w_{0}-w_{0,\rho}^{*}\|=\mathcal{O}(\epsilon)$ as $N=\mathcal{O}(\rho^{-1})$. CA distance takes the order of $\epsilon$ in every iteration by choosing $\rho=\mathcal{O}(\epsilon^{2}),\beta=\mathcal{O}(\epsilon^{4})$, $\alpha=\mathcal{O}(\epsilon^{9})$, and $N=\mathcal{O}(\epsilon^{-2})$. The proof is complete. ∎
E.2 Formal Version and Its Proof of Theorem 4